OpenStudy (anonymous):

GIVE YOU A MEDAL! Which set of equations would be used to solve this word problem?

OpenStudy (anonymous):

An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air.

OpenStudy (anonymous):

An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air let speed in still air be x mph .. speed with head wind = x-40 speed with tail wind = x+40 .. Distance = speed * time

OpenStudy (anonymous):

Headwind distance = 3.5(x-40) Tailwind distance = 2.5(x+40) .. 3.5(x-40) = 2.5(x+40) 3.5x-140=2.5x+100 x=140+100

OpenStudy (anonymous):

add 140 and 1oo and u have ur answer

OpenStudy (anonymous):

x= what mph speed of plane in still air?

OpenStudy (anonymous):

140 + 100 = ???

OpenStudy (anonymous):

d = 3.5(r + 40) and d = 2.5(r – 40) d = 3.5(r – 40) and d = 2.5(r + 40) d = 6(r + 40) None of these systems can be used to solve this problem.

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

yait is d

OpenStudy (anonymous):

thank you can you help me with another?

OpenStudy (anonymous):

sure post it as another questino and and tag me! Dont forget to mark the best response!